OpenStudy (sasogeek):

is -5log(24) the same as -5log(3)-5log(8) or -5log(3)+5log(8) ?

OpenStudy (anonymous):

-5log(3)-5log(8)

OpenStudy (anonymous):

[ @ sasogeek ] - good question. i too want to know ...imformative.

sam (.sam.):

-5log(24) the same as -5log(3)-5log(8) but not -5log(3)+5log(8)

myininaya (myininaya):

\[-5\log(3 \cdot 8)=-5[\log(3)+\log(8)]=-5\log(3)-5\log(8)\]

OpenStudy (anonymous):

I think that's the answer because it's would be: -5[log24] -5[log(3)+log(8)] and you would distribute the -5: so it is: -5long(3)-5log(8)

sam (.sam.):

You can use calculator to check whether the value is same for both too.

OpenStudy (sasogeek):

thanks :)

OpenStudy (anonymous):

np :)

myininaya (myininaya):

To be more general Assume x>0 , y>0 \[a \log(xy)=a [\log(x)+\log(y)] \text{ by one of the properties of loga.} \] \[=a \log(x)+ a \log(y) \text{ by the distributive property}\]